The following information refers to questions 18 and 19

.Assume a U.S.-based MNC is borrowing Romanian leu (ROL) at an interest rate of 8% for one year. Also assume that the spot rate of the leu is $.00012 and the one-year forward rate of the leu is $.00010. The expected spot rate of the leu one-year from now is $.00011.

18) What is the effective financing rate for the MNC assuming it borrows leu on a covered basis?

a) 10%

b) –10%

c) –1%

d) 1%

e) None of the above

19) What is the effective financing rate for the MNC assuming it borrows leu on an uncovered basis?

a) 10%

b) –10%

c) –1%

d) 1%

e) None of the above
Answer #1

18). Effective Rate = [(1 + rROL) * (Forward Rate / Spot Rate)] - 1

= [(1 + 0.08) * (0.00010 / 0.00012)] - 1

= [1.08 * 0.8333] - 1

= 0.9 - 1 = -0.1, or -10%

19). Effective Rate = [(1 + rROL) * Expected Spot Rate] - 1

= [(1 + 0.08) * 0.00011] - 1

= 0.00012 - 1 = -0.9998, or -1%